OpenStudy (anonymous):

2x + y = 7 5x + y = 9 Solve the system of equations.

OpenStudy (anonymous):

so: 5x+y=9 -2x+y=7 ------------ = 3x =2 so x is 1.5 and y is 4

OpenStudy (zpupster):

5(1.5) + 4 = 9 does not work for that equation

OpenStudy (zpupster):

lets try: 5x + y = 9 -2x - y = -7 3x = 2 divide 3 both sides x = 2/3 5(2/3) + y = 9 10/3 + y = 9 y= 9-10/3 y = 17/3 check 2(2/3) +17/3 = 21/3 0r 4/3 + 17/3 = 21/3 or 7......check and 5(2/3) +17/3 = 27/3 or 10/3 + 17/3 = 27/3 or 9..........check

OpenStudy (anonymous):

Thanks for the help :)